# Dense Hopfield Networks in the Teacher-Student Setting

### v1

@@ 3 Teacher-student setting
$\quad+\int dx\frac{1}{\sqrt{2\pi}}\exp\left\{-\frac{1}{2}x^{2}\right\}\Big{\langle}\log\left[\cosh\left(\beta\left[\sqrt{\alpha r}x+\beta^{*}\alpha r^{*}+kz\right]\right)\right]\Big{\rangle}_{z}\Bigg{\}},$
@@
where $z$ is a Rademacher random variable. As in the direct model described in Section 2, the order parameters $m$ and $q$ have a clear interpretation in terms of expected overlaps. $m=\lim_{N\to\infty}\left\langle\frac{1}{N}\sum_{i}\xi_{i}\sigma^{a}_{i}\right\rangle_{\xi^{*},\sigma,\xi}$ is the expected overlap of a retrieval attempt with an example $\sigma^{a}$, and $q=\lim_{N\to\infty}\left\langle\frac{1}{N}\sum_{i}\left\langle\xi_{i}\right\rangle^{2}_{\xi}\right\rangle_{\xi^{*},\sigma}$ is the expected overlap between two retrieval attempts. Similarly, $q^{*}$ is the expected overlap between the teacher and student patterns, i.e. $q^{*}=\lim_{N\to\infty}\left\langle\frac{1}{N}\sum_{i}\xi^{*}_{i}\xi_{i}\right\rangle_{\xi^{*},\sigma,\xi}$. Therefore, it is a good measure of inference performance. The free entropy (Eq. 3) is expected to be exact in absence of mismatch between the teacher and the student, i.e. $\beta^{*}=\beta$. This condition is known as the Nishimori line [45, 46, 47, 48]. Outside of the Nishimori region, RSB corrections are expected. Like in the direct problem, there are different phases characterized by the values of the order parameters:
@@
In sum, when $T^{*}$ is above $T_{\text{crit}}$, the student can only learn the teacher pattern in the signal retrieval phases. In all the other phases, the student pattern is uncorrelated with the signal, being either a random guess ($P$ phase), aligned with a noisy example (inaccurate $eR$ phase), or aligned with a spurious low energy state ($SG$ phase).
@@
In the large-noise scaling, we obtain saddle point equations similar to Eqs. (3) but with $\beta^{*}$ replaced by $\lambda$ (see Appendix D). Conversely, the finite noise scaling leads to
$q^{*}$ $=\Big{\langle}\tanh\left(\beta\left[\eta\alpha r^{*}+kz\right]\right)\Big{\rangle}_{z}$
$m$ $=\Big{\langle}z\tanh\left(\beta\left[\eta\alpha r^{*}+kz\right]\right)\Big{\rangle}_{z}$ (5)
$r^{*}$ $=p\left[q^{*}\right]^{p-1}$
@@
where $\eta$ generally depends on $\beta^{*}$ and $p$ in a non-trivial way, but we find that $\eta=\frac{2\left[\beta^{*}\right]^{2}}{\left(1-2\beta^{*}\right)^{2}}$ when $p=4$ (see Appendix D). These equations can also be derived by extrapolating the large-noise equations equations to $\alpha_{\text{large noise}}\rightarrow 0$ and $\lambda\rightarrow\infty$ with fixed $\lambda\alpha_{\text{large noise}}=\eta\alpha_{\text{finite noise}}$.
@@ 4 Results and Discussion
4.1 Transition to the ordered phases: universality
The paramagnetic solution of Eqs. (3) always exists and is globally stable in the part of the phase diagram where the $T$ is relatively large and $\alpha$ is relatively small. On the other hand, the $gR$ phase exists when $\beta^{2}\alpha p$ and $\beta^{*}\beta\alpha p$ are both large. In fact, in that limit, $q^{*}=q=1$ is a fixed point of Eqs. (3). The critical line where $gR$ becomes globally stable instead of $P$ is not clear from this analysis alone, but we can at least find it analytically in the limit of infinite $p$. As for the direct model, the free entropy and the total entropy of the paramagnetic phase are respectively $\frac{1}{2}\beta^{2}\alpha+\log 2$ and $-\frac{1}{2}\beta^{2}\alpha+\log 2$ [30]. At the same time, the $p\rightarrow\infty$ free entropy takes the form
$f$ $=\operatorname*{Extr}\Bigg{\{}\beta^{*}\beta\alpha\ \theta\left[q^{*}\right]-\frac{1}{2}\beta^{2}\alpha\ \theta\left[q\right]-\beta^{*}\beta\alpha r^{*}q^{*}+\frac{1}{2}\beta^{2}\alpha rq-\frac{1}{2}\beta^{2}\alpha r+\frac{1}{2}\beta^{2}\alpha+\log 2$
$\quad+\int dx\frac{1}{\sqrt{2\pi}}\exp\left\{-\frac{1}{2}x^{2}\right\}\log\left[\cosh\left(\sqrt{\beta^{2}\alpha r}x+\beta^{*}\beta\alpha r^{*}\right)\right]\Bigg{\}},$
where $\theta\left[q\right]:=\lim_{p\rightarrow\infty}q^{p}$, $q\in\left[0,1\right]$, is the Heaviside step function jumping at $q=1$, i.e. $\theta(1)=1$ and $\theta(q)=0$ $\forall q\in[0,1)$. In this limit, the ferromagnetic phase is characterized by $q=q^{*}=1$, and its free entropy is then
@@ 4.1 Transition to the ordered phases: universality
where $T=\sqrt{\frac{\alpha}{2\log 2}}$ is also the temperature where the entropy of the paramagnetic phase becomes negative. Outside of the Nishimori line, this inequality generalizes to $\beta^{*}\beta\alpha>\frac{1}{2}\beta^{2}\alpha+\log 2$, leading to
@@
while the temperature where the paramagnetic free entropy becomes negative stays the same.
In the $p\rightarrow\infty$ limit, the transition towards $gR$ of the inverse model on the Nishimori line is identical to the exact $P$-$SG$ transition of the direct model [30]. We claim that these two critical lines are actually closely related for any $p$. In the Hopfield model with $p=2$, they were already shown to be identical [38]. We will now argue that they overlap for any $p$ and $\beta$ such that the inverse model is outside of the $eR$ phase. In the case of $p=2$, both lines can be obtained exactly from the RS approximation of either the direct model or the inverse model, so there is no obvious advantage to using this equivalence in calculations. In general, while the inverse problem on the Nishimori line is replica symmetric, the direct problem is not, and the $p\geq 3$ replica symmetric $P$-$SG$ transition is not exact. Moreover, even the critical line calculated using 1RSB may be inaccurate due to numerical instability [33]. In this situation, the knowledge of the $gR$ transition in the replica-symmetric inverse problem can be used to locate the exact $P$-$SG$ transition of the direct problem, where symmetry breaking occurs.
For that purpose, we will argue that the direct model is in the paramagnetic phase if and only if the inverse model is in the paramagnetic phase.
@@
As determined in Appendix C and D, the annealed expression $\frac{1}{N}\log\left\langle\mathcal{Z}\right\rangle$ is equal to the free entropy of the paramagnetic phase. Therefore, when the inverse model is in the paramagnetic phase, $P\left(\sigma\right)$ is contiguous to the uniform distribution. This property is called quiet planting and is known to occur more generally in mean-field paramagnets [52, 53, 54, 55]. In our problem setting, it means that if the inverse model is in the paramagnetic phase, then it is equivalent to the direct model. In particular, if the inverse model is in the paramagnetic phase, then so is the direct model. In more intuitive terms, the $gR$ transition temperature of the inverse model must be greater than or equal to the $P$-$SG$ transition temperature of the direct model because the ensemble of examples $\sigma^{a}$ generated by the teacher model is on average at least as structured as the set of random memories stored in the direct model.
@@
This expression is identical to the replicated partition function of the inverse model outside of the $eR$ phase, which therefore must also be in the paramagnetic phase.
As a consequence, the $P$-$SG$ transition line of the direct model must be identical to the $gR$ transition line of the inverse model on the Nishimori line.
@@ 4.2 Phase diagram on the Nishimori line
On the Nishimori line, the student is fully informed about the teacher generative model and uses $\beta=\beta^{*}$ and $p=p^{*}$. In this scenario, thanks to the Nishimori identities [46], it is well known that $\xi^{*}$ and $\xi$ play symmetric roles and that $q^{*}=q$. For the same reason, the overlaps $\frac{1}{N}\sum_{i}\xi^{*}_{i}\xi_{i}$ and $\frac{1}{N}\sum_{i}\xi^{1}_{i}\xi^{2}_{i}$ have the same distribution. From the self-averaging of $\frac{1}{N}\sum_{i}\xi^{*}_{i}\xi_{i}$, it follows that the system is expected to be replica symmetric, and Eqs. (3) and (3) are expected to hold. Fig. (2) shows the phase diagrams obtained by solving the saddle-point equations numerically on the Nishimori line. Both $q^{*}=q$ and the replica symmetry condition are verified. In particular, numerical solutions of a few values of $p\geq 3$ show that the $gR$ transition occurs at a higher $T$ than the line $\beta^{2}\alpha=2\log 2$ where the total entropy of the paramagnetic phase becomes negative. In other terms, the phase transition towards $gR$ prevents the total entropy from becoming negative when $T$ decreases below $\sqrt{\frac{\alpha}{2\log 2}}$, which is consistent with the RS solution being exact on the Nishimori line.
Figure 2: Exact replica-symmetric phase diagrams of the inverse models with $p=3$ on the left, $p=4$ in the center and $p=10$ on the right. Accurate pattern retrieval is not possible in the paramagnetic phase ($P$), but it is possible in the local retrieval phase ($lR$), in the global retrieval phase ($gR$) and in the example retrieval phase ($eR$). The ferromagnetic fixed point corresponding to accurate pattern retrieval is globally stable in the $gR$ phase, but locally stable in the $lR$ phase. The black dashed lines mark the spurious continuation of the $lR$ and $gR$ phase boundaries through the $eR$ phase. The white dashed line is the $p\rightarrow\infty$ $gR$ critical line calculated analytically at the end of Section 3. It matches the corresponding numerical phase boundary increasingly well as $p$ grows larger.
@@
At high $T$, learning is possible only if the amount of examples, i.e. the size of the dataset, is sufficiently large. When $\alpha$ is too small, Eqs. (3) have only a paramagnetic fixed point because the amount of information carried by the dataset is not large enough. Numerical solutions suggest that the paramagnetic fixed point always exist and it is actually locally stable in the whole high-temperature regime. When $\alpha$ is sufficiently large, the signal retrieval fixed point appears as a locally stable attractor. It becomes globally stable as the size of the dataset is increased further.
The critical boundary of the $gR$ phase can be obtained by solving the saddle-point equations numerically (Eqs. 3), and the result is consistent with the analytical $p\rightarrow\infty$ $gR$ boundary of Eq. (6). In fact, we find that the analytical boundary closely agrees with the numerical solution of the saddle-point equations with $p^{*}=p=10$ and remains a good approximation even down to $p^{*}=p=4$.
In the student model, $\sigma$ plays a similar role as the weights of the trainable dense Hopfield network model that Krotov designed for classification of data [26]. In that context, $\xi$ is analogous to the test data whose labels are being predicted (see Fig. 3). In fact, the computation performed by Krotov’s model to recover labels is similar to the update rule used by the student to infer the teacher pattern (see Appendix A). Moreover, the $eR$ and $gR$ phases are respectively reminiscent of the prototype and feature regimes of Krotov’s networks. Therefore, we believe that the student can act as a toy model of label prediction in these two regimes.
@@
Figure 3: The first row of this diagram sketches how a $p$-body Hopfield network in the teacher-student setting can reconstruct an incomplete pattern $\xi^{b}$ to match the teacher pattern $\xi^{*}$ by relying on the examples $\sigma$ obtained from $\xi^{*}$. The second row summarizes how a dense neural network trained by Krotov can recover the labels $y^{\prime}$ of the data $x$ given the weights $w$ learned from $x$ [26]. Both models tackle similar tasks using an approach where $\sigma$ and $\xi^{b}$ respectively play the same roles as $w$ and $(x,y^{\prime})$. The forward propagation algorithm used to generate $y^{\prime}$ is similar to the update rule of the student (see [26] and Appendix A), but the backpropagation algorithm used to learn $w$ is very different from the update rule of the teacher.
@@
Fig. (4) shows results from Monte Carlo simulations with $p=3$, where $L$ replicas of the student pattern $\{\xi^{b}\}_{b=1}^{L}$ are initialized to the teacher pattern $\xi^{*}$ corrupted by some Rademacher noise $\varepsilon$. In other words, the initial values of $\xi^{b}_{i}$ are sampled from $\left(1-\varepsilon\right)\delta\left(\xi_{i}-\xi^{*}_{i}\right)+\frac{\varepsilon}{2}\left[\delta\left(\xi_{i}+1\right)+\delta\left(\xi_{i}-1\right)\right]$ with $\varepsilon\in[0,1]$. The value of $\varepsilon$ is tuned so that the simulations start relatively close to the saddle-point solutions. As explained previously, $gR$ is a hard phase, so this initialization is necessary to make $\xi^{b}$ converge to $gR$ in a reasonable amount of time. Additionally, it is also used to make $\xi^{b}$ converge to the $lR$ phase rather than the $P$ phase when desired. Once the simulations are over, the overlaps are averaged over all $L$ replicas. If we fix $\varepsilon=0$, then the simulations generally converge to the $lR$ phase when it is a fixed point. If instead we initialize them to the saddle-point solutions by handpicking $\varepsilon$, then they stay near the initial overlaps. In either case, the simulations converge to $eR$ when it is globally stable. Some simulation data points might be systematically shifted up with respect to the saddle-point solutions. However, this difference decreases with the system size $N$, so finite size effects seem sufficient to explain it (see Fig. 9 in Appendix F). Overall, the Monte-Carlo simulations are in very good agreement with the $p=3$ overlap landscape obtained by solving the saddle-point equations numerically.

### v2

@@ 3.1 Matched interaction orders
$\quad+\int dx\frac{1}{\sqrt{2\pi}}\exp\left\{-\frac{1}{2}x^{2}\right\}\Big{\langle}\log\left[\cosh\left(\beta\left[\sqrt{\alpha r}x+\beta^{*}\alpha r^{*}+kz\right]\right)\right]\Big{\rangle}_{z}\Bigg{\}}\,,$ (4)
@@
where $z$ is a Rademacher random variable and $\alpha=\frac{Mp!}{N^{p-1}}$. As in the direct model described in Section 2, the order parameters $m$ and $q$ have a clear interpretation in terms of expected overlaps. $m=\lim_{N\to\infty}\left\langle\frac{1}{N}\sum_{i}\xi_{i}\sigma^{a}_{i}\right\rangle_{\xi^{*},\sigma,\xi}$ is the expected overlap of a retrieval attempt with an example $\sigma^{a}$, and $q=\lim_{N\to\infty}\left\langle\frac{1}{N}\sum_{i}\left\langle\xi_{i}\right\rangle^{2}_{\xi}\right\rangle_{\xi^{*},\sigma}$ is the expected overlap between two retrieval attempts. Similarly, $q^{*}$ is the expected overlap between the teacher and student patterns, i.e. $q^{*}=\lim_{N\to\infty}\left\langle\frac{1}{N}\sum_{i}\xi^{*}_{i}\xi_{i}\right\rangle_{\xi^{*},\sigma,\xi}$. Therefore, it is a good measure of inference performance. The free entropy (Eq. 3.1) is expected to be exact in absence of mismatch between the teacher and the student, i.e. $\beta^{*}=\beta$. This condition is known as the Nishimori line [45, 46, 47, 48]. Outside of the Nishimori region, RSB corrections are expected. Like the direct problem, the inverse problem with $T^{*}>T_{\text{crit}}$ has different phases characterized by the values of the order parameters:
@@
In sum, when $T^{*}$ is above $T_{\text{crit}}$, the student can only learn the teacher pattern in the signal retrieval phases. In all the other phases, the student pattern is uncorrelated with the signal, being either a random guess ($P$ phase), aligned with a noisy example (inaccurate $eR$ phase), or aligned with a spurious low energy state ($SG$ phase). We stress that we cannot have $m\neq 0$ and $q^{*}\neq 0$ at the same time (accurate $eR$ phase) when $T^{*}>T_{\text{crit}}$ because $\lim_{N\to\infty}\left\langle\frac{1}{N}\sum_{i}\xi^{*}_{i}\sigma^{a}_{i}\right\rangle_{\xi^{*},\sigma}=0$ in that regime (see Fig. 1, $\alpha=0$ axis).
3.2 Mismatched interaction orders
@@
In the large-noise scaling, we obtain saddle point equations similar to Eqs. (3.1) but with $\beta^{*}$ replaced by $\lambda$ (see Appendix D). Conversely, the finite noise scaling leads to
$q^{*}$ $=\Big{\langle}\tanh\left(\beta\left[\eta\alpha r^{*}+kz\right]\right)\Big{\rangle}_{z}\,,$
$m$ $=\Big{\langle}z\tanh\left(\beta\left[\eta\alpha r^{*}+kz\right]\right)\Big{\rangle}_{z}\,,$ (6)
$r^{*}$ $=p\left[q^{*}\right]^{p-1}\,,$
@@
where $\eta$ generally depends on $\beta^{*}$ and $p$ in a non-trivial way, but we find that $\eta=\frac{2\left[\beta^{*}\right]^{2}}{\left(1-2\beta^{*}\right)^{2}}$ when $p=4$ (see Appendix D). These equations can also be derived by extrapolating the large-noise equations to $\alpha_{\text{large noise}}\rightarrow 0$ and $\lambda\rightarrow\infty$ with fixed $\lambda\alpha_{\text{large noise}}=\eta\alpha_{\text{finite noise}}$.
@@ 4 Results and Discussion
4.1 Retrieval transition at large interaction order
The paramagnetic solution of Eqs. (3.1) always exists and is globally stable in the part of the phase diagram where the temperature $T$ is relatively large and $\alpha=\frac{Mp!}{N^{p-1}}$ is relatively small. On the other hand, the $gR$ phase exists when $\beta^{2}\alpha p$ and $\beta^{*}\beta\alpha p$ are both large. In fact, in that limit, $q^{*}=q=1$ is a fixed point of Eqs. (3.1). The critical line where $gR$ becomes globally stable instead of $P$ is not clear from this analysis alone, but we can at least find it analytically in the limit of infinite $p$. As for the direct model, the free entropy and the total entropy of the paramagnetic phase are respectively $\frac{1}{2}\beta^{2}\alpha+\log 2$ and $-\frac{1}{2}\beta^{2}\alpha+\log 2$ [30]. At the same time, the $p\rightarrow\infty$ free entropy takes the form
$f$ $=\operatorname*{Extr}\Bigg{\{}\beta^{*}\beta\alpha\ \theta\left(q^{*}-1\right)-\frac{1}{2}\beta^{2}\alpha\ \theta\left(q-1\right)-\beta^{*}\beta\alpha r^{*}q^{*}+\frac{1}{2}\beta^{2}\alpha rq-\frac{1}{2}\beta^{2}\alpha r+\frac{1}{2}\beta^{2}\alpha$
$\quad+\log 2+\int dx\frac{1}{\sqrt{2\pi}}\exp\left\{-\frac{1}{2}x^{2}\right\}\log\left[\cosh\left(\sqrt{\beta^{2}\alpha r}x+\beta^{*}\beta\alpha r^{*}\right)\right]\Bigg{\}}\,,$
where $\theta\left(q-1\right):=\lim_{p\rightarrow\infty}q^{p}$, $q\in\left[0,1\right]$, is the Heaviside step function jumping at $q=1$, i.e. $\theta(1)=1$ and $\theta(q)=0$ $\forall q\in[0,1)$. In this limit, the ferromagnetic phase is characterized by $q=q^{*}=1$, and its free entropy is then
@@ 4.1 Retrieval transition at large interaction order
where $T_{E}=\sqrt{\frac{\alpha}{2\log 2}}$ is also the temperature below which the total entropy of the paramagnetic phase becomes negative. Outside of the Nishimori line, this inequality generalizes to $\beta^{*}\beta\alpha>\frac{1}{2}\beta^{2}\alpha+\log 2$, leading to
@@
while the temperature where the paramagnetic total entropy becomes negative stays the same.
4.2 Transition to the ordered phases: Universality
In the $p\rightarrow\infty$ limit, the transition towards $gR$ of the inverse model on the Nishimori line is identical to the exact $P$-$SG$ transition of the direct model [30]. We claim that these two critical lines are actually closely related for any $p$. In the Hopfield model with $p=2$, they were already shown to be identical [38]. We will now argue that they overlap for any $p$ and $\beta$ such that $T>T_{\text{crit}}$ (see Figs. 2 and 1). In the case of $p=2$, both lines can be obtained exactly from the RS approximation of either the direct model or the inverse model, so there is no obvious advantage to using this equivalence in calculations. In general, while the inverse problem on the Nishimori line is replica symmetric, the direct problem is not, and the $p\geq 3$ replica symmetric $P$-$SG$ transition is not exact. Moreover, even the critical line calculated using 1RSB may be inaccurate due to numerical instability [33]. In this situation, the knowledge of the $gR$ transition in the replica-symmetric inverse problem can be used to locate the exact $P$-$SG$ transition of the direct problem, where symmetry breaking occurs.
For that purpose, we will argue that, given $T>T_{\text{crit}}$, the direct model is in the paramagnetic phase if and only if the inverse model is in the paramagnetic phase.
@@ 4.2 Transition to the ordered phases: Universality
As determined in Appendix C and D, the annealed expression $\frac{1}{N}\log\left\langle\mathcal{Z}\right\rangle$ is equal to the free entropy of the paramagnetic phase. Therefore, when the inverse model is in the paramagnetic phase, $P\left(\sigma\right)$ is contiguous to the uniform distribution. This property is called quiet planting and is known to occur more generally in mean-field paramagnets [58, 59, 60, 61]. In our problem setting, it means that if the inverse model is in the paramagnetic phase, then it is equivalent to the direct model. In particular, if the inverse model is in the paramagnetic phase, then so is the direct model. In more intuitive terms, the $gR$ transition temperature of the inverse model must be greater than or equal to the $P$-$SG$ transition temperature of the direct model because the ensemble of examples $\sigma^{a}$ generated by the teacher model is on average at least as structured as the set of i.i.d. random memories stored in the direct model.
@@
This expression is identical to the replicated partition function of the inverse model with $T>T_{\text{crit}}$, which therefore must also be in the paramagnetic phase.
As a consequence, when $T>T_{\text{crit}}$, the $P$-$SG$ transition line of the direct model must be identical to the $gR$ transition line of the inverse model on the Nishimori line.
@@ 4.3 Phase diagram on the Nishimori line
On the Nishimori line, the student is fully informed about the teacher generative model and uses $\beta=\beta^{*}$ and $p=p^{*}$. In this scenario, thanks to the Nishimori identities [46], it is well known that $\xi^{*}$ and $\xi$ play symmetric roles and that $q^{*}=q$. For the same reason, the overlaps $\frac{1}{N}\sum_{i}\xi^{*}_{i}\xi_{i}$ and $\frac{1}{N}\sum_{i}\xi^{1}_{i}\xi^{2}_{i}$ have the same distribution. From the self-averaging of $\frac{1}{N}\sum_{i}\xi^{*}_{i}\xi_{i}$, it follows that the system is expected to be replica symmetric, and Eqs. (3.1) and (3.1) are expected to hold. Fig. (2) shows the phase diagrams obtained by solving the saddle-point equations numerically on the Nishimori line. Both $q^{*}=q$ and the replica symmetry condition are verified. In particular, numerical solutions of a few values of $p\geq 3$ show that the $gR$ transition occurs at a higher $T$ than the line $\beta^{2}\alpha=2\log 2$ where the total entropy of the paramagnetic phase becomes negative. In other terms, the phase transition towards $gR$ prevents the total entropy from becoming negative when $T$ decreases below $\sqrt{\frac{\alpha}{2\log 2}}$, which is consistent with the RS solution being exact on the Nishimori line.
Figure 2: Exact RS phase diagrams of inverse models on the Nishimori line, i.e. $p^{*}=p$ and $\beta^{*}=\beta$. the left, center and right plots respectively have $p=3$, $p=4$ and $p=10$. Accurate pattern retrieval is not possible in the paramagnetic phase ($P$), but it is possible in the local retrieval phase ($lR$), in the global retrieval phase ($gR$) and in the example retrieval phase ($eR$). The ferromagnetic fixed point corresponding to accurate pattern retrieval is globally stable in the $gR$ phase, but locally stable in the $lR$ phase. The critical temperature of the $eR$ phase is the critical temperature $T_{\text{crit}}$ of the direct problem with one pattern (see Fig. 1, $\alpha=0$ axis). The black dashed lines mark the spurious continuation of the $lR$ and $gR$ phase boundaries through the $eR$ phase. The white dashed line is the $p\rightarrow\infty$ $gR$ critical line calculated analytically in Section 4.1. It matches the corresponding numerical phase boundary increasingly well as $p$ grows larger. The white dotted lines on the $p=3$ plot mark the 1RSB and d1RSB critical temperatures $T_{s}\left(\alpha,3\right)$ and $T_{d}\left(\alpha,3\right)$ of the direct model (see Section 2). We truncated them below $T_{\text{crit}}$ for improved visibility. $T_{s}\left(\alpha,3\right)$ and $T_{d}\left(\alpha,3\right)$ are obtained by rescaling the corresponding critical temperatures found in [54] by $\sqrt{2\alpha}$.
@@
At high $T$, learning is possible only if the amount of examples, i.e. the size of the dataset, is sufficiently large. When $\alpha$ is too small, Eqs. (3.1) have only a paramagnetic fixed point because the amount of information carried by the dataset is not large enough. Numerical solutions suggest that the paramagnetic fixed point always exist and it is actually locally stable in the whole high-temperature regime. When $\alpha$ is sufficiently large, the signal retrieval fixed point appears as a locally stable attractor ($lR$ phase). It becomes globally stable ($gR$ phase) as the size of the dataset is increased further or the student temperature decreases.
As per the previous Section, the critical boundary of the $gR$ phase obtained by solving Eqs. 3.1 is identical to the 1RSB $P$-$SG$ transition temperature $T_{s}(\alpha,p)$ of the direct model. Similarly, we observe that the metastable $lR$ phase coincides with the d1RSB phase of the direct model (see Fig. 2). Our results are also consistent with the fact that $T_{s}(\alpha,p)\to T_{E}(\alpha)$ in the $p\rightarrow\infty$ limit. In fact, we find that the analytical limit boundary closely agrees with the numerical solution of the saddle-point equations with $p^{*}=p=10$ and remains a good approximation even down to $p^{*}=p=4$.
In the student model, $\sigma$ plays a similar role as the weights of the trainable dense Hopfield network model that K & H designed for classification of data [26]. In that context, $\xi$ is analogous to the test data whose labels are being predicted (see Fig. 3). In fact, the computation performed by K & H’s model to recover labels is similar to the update rule used by the student to infer the teacher pattern (see Appendix A). Moreover, the $eR$ and $gR$ phases are respectively reminiscent of the prototype and feature regimes of K & H’s networks. Therefore, we believe that the student can act as a toy model of label prediction in these two regimes.
@@
Figure 3: The first row of this diagram sketches how a $p$-body Hopfield network in the teacher-student setting can reconstruct an incomplete pattern $\xi^{b}$ to match the teacher pattern $\xi^{*}$ by relying on the examples $\sigma$ obtained from $\xi^{*}$. The second row summarizes how a dense neural network trained by K & H can recover the labels $y^{\prime}$ of the data $x$ given the weights $w$ learned from $x$ [26]. Both models tackle similar tasks using an approach where $\sigma$ and $\xi^{b}$ respectively play the same roles as $w$ and $(x,y^{\prime})$. The forward propagation algorithm used to generate $y^{\prime}$ is similar to the update rule of the student (see [26] and Appendix A), but the backpropagation algorithm used to learn $w$ is very different from the update rule of the teacher.
@@
Fig. (4) shows results from Monte Carlo simulations with $p=3$, where $L$ replicas of the student pattern $\{\xi^{b}\}_{b=1}^{L}$ are initialized to the teacher pattern $\xi^{*}$ corrupted by some Rademacher noise $\varepsilon$. In other words, the initial values of $\xi^{b}_{i}$ are sampled from the distribution $\left(1-\varepsilon\right)\delta\left(\xi_{i}-\xi^{*}_{i}\right)+\frac{\varepsilon}{2}\left[\delta\left(\xi_{i}+1\right)+\delta\left(\xi_{i}-1\right)\right]$ with $\varepsilon\in[0,1]$. The value of $\varepsilon$ is tuned so that the simulations start relatively close to the saddle-point solutions. As explained previously, $gR$ is a hard phase, so this initialization is necessary to make $\xi^{b}$ converge to $gR$ in a reasonable amount of time. Additionally, it is also used to make $\xi^{b}$ converge to the $lR$ phase rather than the $P$ phase when desired. Once the simulations are over, the overlaps are averaged over all $L$ replicas. If we fix $\varepsilon=0$, then the simulations generally converge to the $lR$ phase when it is a fixed point. If instead we initialize them to the saddle-point solutions by handpicking $\varepsilon$, then they stay near the initial overlaps. In either case, the simulations converge to $eR$ when it is globally stable. Some simulation data points might be systematically shifted up with respect to the saddle-point solutions. However, this difference decreases with the system size $N$, so finite size effects seem sufficient to explain it (see Fig. 9 in Appendix F). Overall, the Monte-Carlo simulations are in very good agreement with the $p=3$ overlap landscape obtained by solving the saddle-point equations numerically.